OpenStudy (anonymous):

Let f(x)= –2x + 4 and g(x) = –6x – 7. Find f(x)– g(x).

OpenStudy (saifoo.khan):

Since we have to find f(x) - g(x) Write it this way: f(x) - g(x) (–2x + 4) - (–6x – 7) Simplify.

OpenStudy (anonymous):

how do you simplify?

OpenStudy (saifoo.khan):

Add like terms.

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

what you would do is -2x+-6x to get? and 4+-7 to get? And you'll have your answer

OpenStudy (anonymous):

wait what? i'm so confused..

OpenStudy (anonymous):

-2x and -6x are like terms because of the X's so add those.. and 4 and -7 are alike as well because they don't have an X. Add those

OpenStudy (phi):

How far did you get? Do you understand this step f(x) - g(x)= (–2x + 4) - (–6x – 7)

OpenStudy (anonymous):

nope

OpenStudy (phi):

The question is f(x)= –2x + 4 and g(x) = –6x – 7. Find f(x)– g(x). f(x) is short-hand for -2x+4 (think of it as a nickname) g(x) is short-hand for –6x – 7 when someone wants f(x)-g(x) they want (-2x+4)- (-5x-7) we say f(x) - g(x)= (–2x + 4) - (–6x – 7) so there is not too much to understand, except that you can subtract complicated things like (-2x+4) and (-6x-7)

OpenStudy (phi):

*typo in when someone wants f(x)-g(x) they want (-2x+4)- (-5x-7) should be (-2x+4)- (-6x-7)

OpenStudy (phi):

clear as mud?

OpenStudy (phi):

but you are not finished because you can simplify this f(x) - g(x)= (–2x + 4) - (–6x – 7)

OpenStudy (phi):

f(x) - g(x)= (–2x + 4) - (–6x – 7) we should write this as f(x) - g(x)= (–2x + 4) + -1* (–6x – 7) (subtraction is the same as adding a negative number, and a negative number is -1 times that number) the first step is "distribute" the minus one in -1*(–6x – 7) this is a "rule" you should learn: distribute means multiply the -1 times each term inside can you do that?

OpenStudy (anonymous):

i'll try..

OpenStudy (anonymous):

so it's 2x +4 -6 -7?

OpenStudy (phi):

Let's concentrate on -1*(–6x – 7) it is best to think of this as -1* (-6*x + -7) btw, remember that -6x means -6*x (-6 times x) distribute the -1 means multiply -1 times each term inside the parens -1* -6*x + -1*-7 now we can simplify this. what do we get ?

OpenStudy (phi):

if it were -1*-6* 8 you could do it, right? what about doing it half-way: -1*-6 is +6 so we get 6*8 if we had x instead it would be 6x -1*-6*x is 6x

OpenStudy (anonymous):

okay so i's 6x+8?

OpenStudy (phi):

almost. what is -1*(–6x – 7) ?

OpenStudy (phi):

I should write is as -1*(–6x + –7)

OpenStudy (anonymous):

6x+8

OpenStudy (phi):

so you are saying -1*-7 is 8?

OpenStudy (anonymous):

isn't it?

OpenStudy (phi):

1+7 is 8 -1+ -7 is -8 by 1*7 is 7 and -1* 7 is -7 and 1* -7 is -7 what is -1* -7 ?

OpenStudy (anonymous):

7?

OpenStudy (phi):

yes. so what is -1*(–6x + –7)

OpenStudy (anonymous):

6x+7?

OpenStudy (phi):

yes. so far we have f(x) - g(x)= (–2x + 4) - (–6x – 7) becomes f(x) - g(x)= (–2x + 4) +6x+7

OpenStudy (phi):

we do not need the parentheses, so we could write this as f(x) - g(x)= –2x + 4 +6x+7 you know we can add numbers in any order so we could write this as f(x) - g(x)= 6x + -2x +4 + 7

OpenStudy (phi):

now you can do some of that, right? you can add 4 and 7 and put down one number instead of 4+7

OpenStudy (anonymous):

8x+11?

OpenStudy (phi):

the 11 is correct. 4+7 is 11 you should get used to writing the whole equation (good practice) f(x) - g(x)= 6x + -2x +11 now for the x's you have 6 x's and take away 2 x's how many x's do you have?

OpenStudy (anonymous):

4

OpenStudy (phi):

yes you have 4 x's we write that as 4x so finally, what is f(x) - g(x)= 6x + -2x +11 write down the whole equation

OpenStudy (anonymous):

4x+11?

OpenStudy (phi):

yes, but that isn't the whole equation.

OpenStudy (anonymous):

so what is it?

OpenStudy (phi):

you should write the left hand side, the = sign, and then the right-hand side

OpenStudy (anonymous):

okay now i'm confused again

OpenStudy (phi):

like this f(x) - g(x)= 4x+11

OpenStudy (anonymous):

oh okay, so that's my answer then?

OpenStudy (phi):

yes

OpenStudy (anonymous):

okay thank you for your help! :)